Alice is a waitress and collects her tips at the table. At the end of the shift she has 68 bills in her tip wallet, all ones and fi ves. If the total value of her tips is 172, how many of each bill does she have?
Answer by Cromlix(4381)

$1 = x
$5 = y
x + y = 68 ...1
x + 5y = 172....2
Multiply (1) by 5
5x + 5y = 340...3
x + 5y = 172....2
Subtract (2) from (3)
4x = 168
x = 42
Substitute x = 42 into (1)
x + y = 68
42 + y = 68
y = 68 - 42
y = 26
Alice had 42 x $1 and 26 x $5.
Hope this helps.
:-)